OpenStudy (anonymous):

Solve √(5)^(8n)= 125^(n + 3)

OpenStudy (anonymous):

Express both sides in terms of powers of 5

OpenStudy (anonymous):

equate the powers

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

8n=n+3 7n=3 n=7/3

OpenStudy (anonymous):

125 = 5^3 sqrt of 5 = 5^(1/2) can you do it now?

OpenStudy (anonymous):

5^x = 5^y => x = y that's the logic